OpenStudy (anonymous):

simplify the expression 36^1/2

OpenStudy (anonymous):

help me cuties :)

OpenStudy (anonymous):

18.

OpenStudy (anonymous):

thank you

Parth (parthkohli):

I'd help if you don't use the word "cuties" for us. lol \[a^{\frac{1}{2}} = \sqrt{a}\]@Jada<3 careful...

OpenStudy (anonymous):

Answer = 6

OpenStudy (anonymous):

I Google It .. Kmsl ! Butt Ahhh Yea

OpenStudy (anonymous):

36^1/2 = square root of 36. Where 6 is the answer.

OpenStudy (anonymous):

what about